1. Suppose the pleasure Romeo gets from giving the fair Juliet roses for valentines day next weekend can be described by a dollar benefit function B(r)= 3.4r - 0.1r^2

a. What is the equation of romeos marginal benefits, MB(r)?

b. What is romeos marginal benefit if he gave Juliet 10 roses?

c. Suppose roses cost $1 each. How many boded will Romeo give to Juliet?

Show how you got your answers.
Answer #1

A) B(r)=3.4r-0.1*r^2

MB= DERIVATIVE of total Benefit

MB=3.4-0.2r

B) r=10

MB=3.4-0.2*10=3.4-2=1.4

C) pr=$1

The optimal number of roses is where,

MB=pr

So,

3.4-0.2r=1

2.4=0.2r

r=2.4/0.2=12

So Romeo will give 12 roses to Juliet